OpenStudy (anonymous):

Choose 1 system of equations from the choices below. Then, solve the system and post your solution, showing your steps so that other students can see which method you chose.

OpenStudy (anonymous):

–y + 3x = 6 y = –6x + 12 6x – 4y = 54 –9x + 2y = –69 2y = x + 1 –2x – y = 7

OpenStudy (anonymous):

Choose one!

OpenStudy (anonymous):

–y + 3x = 6 y = –6x + 12

OpenStudy (anonymous):

@chaguanas

OpenStudy (anonymous):

Step 2: solve the solution!